In two years time jam will be as old as dindo is now. in 26 years dindo will be twica as old as jam is now . how old is dindo?
Answer by Theo(13342)

let x = jam's age now.
let y = dindo's age now.
in 2 years jam will be as old as dindo is now.
this means that x + 2 = y
in 26 years dindo will be twice as old as jam is now.
this means that y + 26 = 2 * x
you have 2 equations that need to be solved simultaneously.
the first equation is:
y = x + 2
the second equation is:
y + 26 = 2 * x
the trick in solving 2 equations in 2 unknowns is to reduce the number of unknowns and equations to 1 which can then be solved. after you solve for the 1 unknown, then you can solve for the other.
in this problem, we'll use substitution as the method to reduce the 2 equations in 2 unknowns to 1 equation in 1 unknown.
the first equation says that y = x + 2.
we'll substitute for y in the second equation to get:
y + 26 = 2 * x becomes:
(x + 2) + 26 = 2 * x
that last equation is now 1 equation in 1 unknown.
simplify it to get:
x + 28 = 2 * x
subtract x from both sides of the equation to get:
28 = x
this is the same as:
x = 28
you now know what x is and you can solve for y.
use the first equation to get:
y = x + 2 which becomes:
y = 30
you now should have both x and y and you can confirm them as being good or not good by substituting in the original equations.
the first equation is y = x + 2
since y = 30 and x = 2, this equation is true and the values of x and y are confirmed as good.
the second equation is y + 26 = 2 * x
y + 26 = 30 + 26 = 56 and 2 * x = 2 * 28 = 56, so the equation is true and the value of x and y are confirmed as being good for the second equation.
since the value of x and y are confirmed as good in both equations, then the value of x and y have simultaneously solved both equations.
the answer is:
x = 28
y = 30
this means that jam is 28 and dindo is 30.
in 2 years jam will be 30 which is how old dindo is now.
in 26 years dindo will be 30 + 26 = 56 years old which is 2 * how old jam is now (jam is now 28 years old).